OpenStudy (anonymous):

The integral of 5xsqrt(1-x^4) = ?

OpenStudy (anonymous):

\[\int\limits_{}^{}5x \sqrt{1-x^4}dx\]

OpenStudy (anonymous):

Please help. Ive tried Trig Sub but im doing something incorrect

OpenStudy (anonymous):

Hmmm, not completely sure myself.

OpenStudy (anonymous):

What if you let \[ x^2=\sin(\theta) \]

OpenStudy (anonymous):

I'll show you what I attempted.

OpenStudy (anonymous):

\[\int\limits_{}^{}5x \sqrt{1-x^4}dx\] \[5\int\limits_{}^{}x \sqrt{1-x^4}dx\] \[5\int\limits_{}^{}x \sqrt{1-(x^2)^2}dx\] \[5\int\limits_{}^{}x \sqrt{1-\sin^2(\theta})^2dx\]

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

Do you know how to do a substitution such as \(x^2=\sin(\theta)\)?

OpenStudy (anonymous):

Not sure.

OpenStudy (anonymous):

if x squared = sin(theta) the radical would be 1-sin^2(theta), right?

OpenStudy (anonymous):

Differentiate both sides with respect to \(\theta\) to get: \[ 2x\frac{dx}{d\theta}=\cos(\theta) \]

OpenStudy (anonymous):

so \[1/2(5)\int\limits_{}^{}(\sqrt{\sin(\theta)}\sqrt{1-\sin^2(\theta)} )d( \theta)\]

OpenStudy (anonymous):

We can simplify this: \[ \frac{dx}{d\theta} = \frac{\cos(\theta)}{2x}\implies dx=\frac{\cos(\theta)}{2x}d\theta \implies 2x\;dx=\cos(\theta) \;d\theta \]

OpenStudy (anonymous):

So \[ \frac{5}{2}\int \sqrt{1-(x^2)^2}\cdot 2x\;dx =\frac{5}{2}\int\sqrt{1-\sin^2(\theta)}\;d\theta \]

OpenStudy (anonymous):

So \[5/2\int\limits_{}^{}(\sqrt{\cos^2(\theta)})d(\theta)\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[5/2\int\limits_{}^{}\cos(\theta)d(\theta)\]

OpenStudy (anonymous):

Should be \(|\cos(\theta)|\)

OpenStudy (anonymous):

\[5/2(\sin(\theta))+C\]

OpenStudy (anonymous):

so \[5/2x^2+C\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Ummm, does it make sense when you differentiate it?

OpenStudy (anonymous):

It doesnt make sense and Im confused.

OpenStudy (anonymous):

Hmm, if we did \(x=\sqrt{\sin(\theta)}\) maybe...

OpenStudy (anonymous):

I'm not sure, let me think.

OpenStudy (anonymous):

I appreciate it.

OpenStudy (anonymous):

Let's start with \(u=x^2\) so \(du=2xdx\)\[ \frac{5}{2}\int\sqrt{1-(x^2)^2}2xdx = \frac{5}{2}\int\sqrt{1-u^2}du \]

OpenStudy (anonymous):

Then we do \(u = \sin\theta\) so \(du=\cos\theta d\theta\)\[ \frac{5}{2}\int\sqrt{1-u^2}du=\frac 52\int \sqrt{1-\sin^2\theta}\cos\theta d\theta \]

OpenStudy (anonymous):

That gives us \[ \frac52\int \cos^3\theta d\theta \]

OpenStudy (anonymous):

Right, Im following.

OpenStudy (anonymous):

I should say \[ \frac 52\int \cos^2\theta d\theta \]

OpenStudy (anonymous):

then \[ \frac54\int 1+\cos(2\theta)d\theta \]

OpenStudy (anonymous):

Ohh, I see.

OpenStudy (anonymous):

I found the error with my previous substitution. I put: \[ \frac{5}{2}\int \sqrt{1-(x^2)^2}\cdot 2x\;dx =\frac{5}{2}\int\sqrt{1-\sin^2(\theta)}\;d\theta \]It should have been\[ \frac{5}{2}\int \sqrt{1-(x^2)^2}\cdot 2x\;dx =\frac{5}{2}\int\sqrt{1-\sin^2(\theta)}\;\color{red}{\cos(\theta)}d\theta \]

OpenStudy (anonymous):

How did we not see that? You're right. So let u = sin theta and du = cos theta. Then 5/2 integral of rad (1-u^2)du ?

OpenStudy (anonymous):

Well \(x^2=\sin(\theta)\) or \(x^2=u=\sin(\theta)\) both lead to the same result.

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Just continue where I left off.

OpenStudy (anonymous):

so 5/2 integral of (cos theta)^2

OpenStudy (anonymous):

Use half angle.

OpenStudy (anonymous):

(1+cos(2 theta))/2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you very much @wio